# Personalized Pieces: Efficient Personalized Large Language Models through Collaborative Efforts

### v1

@@ Abstract
Personalized large language models (LLMs) aim to tailor interactions, content, and recommendations to individual user preferences. While parameter-efficient fine-tuning (PEFT) methods excel in performance and generalization, they are costly and limit communal benefits when used individually. To this end, we introduce Personalized Pieces (Per-Pcs), a framework that allows users to safely share and assemble personalized PEFT efficiently with collaborative efforts. Per-Pcs involves selecting sharers, breaking their PEFT into pieces, and training gates for each piece. These pieces are added to a pool, from which target users can select and assemble personalized PEFT using their history data. This approach preserves privacy and enables fine-grained user modeling without excessive storage and computation demands. Experimental results show Per-Pcs outperforms non-personalized and PEFT retrieval baselines, offering performance comparable to OPPU with significantly lower resource use across six tasks. Further analysis highlights Per-Pcs’s robustness concerning sharer count and selection strategy, pieces sharing ratio, and scalability in computation time and storage space. Per-Pcs’s modularity promotes safe sharing, making LLM personalization more efficient, effective, and widely accessible through collaborative efforts.111Code will be available at https://github.com/TamSiuhin/Per-Pcs
@@
Zhaoxuan Tan, Zheyuan Liu, Meng Jiang University of Notre Dame {ztan3, zliu29, mjiang2}@nd.edu
@@ 3 Experiment Settings
We adopt the Large Language Model Personalization (LaMP) benchmark Salemi et al. (2023) for our experiments, which consists of six public language model personalization tasks, including three text classification tasks (personalized citation identification, movie tagging, and producing rating) and three text generation tasks (personalized news headline generation, scholarly title generation, and tweet paraphrasing).333Task details can be found in Appendix F. We exclude the LaMP-6: Email subject generation task since it involves private data that we cannot access. We randomly select 25% of users to train the base model for task adaptation. From the remaining users, we randomly sample 100 to serve as test users for efficient and fair comparison with OPPU Tan et al. (2024). The rest of the users are used as sharer candidates who consent to share their PEFT parameters.444Statistics are presented in Table 4.
@@
We compare our proposed Per-Pcs with the non-personalized baseline, prompt-based methods (retrieval-augmented Salemi et al. (2023) and profile-augmented personalization Richardson et al. (2023)), and PEFT-based personalization methods (PEFT retrieval Zhao et al. (2024) and OPPU Tan et al. (2024)). Although PEFT retrieval has not been applied to personalization before, we employ it as a PEFT-level composition baseline. compared with Per-Pcs, OPPU requires significantly more resources, which can be seen as the upper bound for sharer personal PEFT composition. We provide more baseline details in Appendix G. For all baselines and Per-Pcs, we use Llama-2-7B Touvron et al. (2023) as the base LLM and BM25 Trotman et al. (2014) for retrieval operations to ensure efficient and fair comparisons.
@@ 6.2 Model Parameter Composition
Existing work has shown that performing weighted linear interpolation of model parameters leads to the composition of each model ability Li et al. (2022); Tam et al. (2023). This approach recycles efforts and computational resources used to create specialized models. These methods can be divided into model-, PEFT-, and piece-level compositions.
Model-level composition methods treat the entire model parameter as the minimum composition unit Wortsman et al. (2022); Choshen et al. (2022); Ramé et al. (2023); Jin et al. (2022). Ilharco et al. (2022) propose the task vector, which subtracts the weights of a fine-tuned model from the pre-trained weights and conducts task vector arithmetic to enable generalization across tasks and domains. PEFT offers lightweight alternatives for fine-tuning LLMs by updating small, plug-in parameters while keeping the pre-trained weights frozen to save computational resources He et al. (2021). In PEFT-level composition, the entire PEFT module is treated as the minimum unit. By composing PEFT parameters, models can achieve task and domain generalization Shah et al. (2023); Gou et al. (2023); Zhang et al. (2023). LoRAHub Huang et al. (2023) uses a black-box optimizer to integrate specialized LoRAs, facilitating generalization to unseen tasks. Another line of work focuses on retrieving PEFT Jang et al. (2023b). LoRARetriever Zhao et al. (2024) retrieves and composes multiple LoRAs based on the given input. In piece-level composition, the minimum composition unit is a plug-in sub-component of PEFT within a specific layer. For instance, in LoRA, each low-rank update at a linear layer constitutes a "piece." Muqeeth et al. (2024) focuses on task generalization and proposes recycling PEFT pieces by employing per-token and per-piece composition under zero-shot settings.
@@ 9 Ethical Considerations
Appendix A Ablation Study
Appendix B Computation Resources Details
@@ Appendix C Hyperparameters
Appendix D Modeling Users with Different Active Levels
Appendix E Scientific Artifacts
Per-Pcs is built with the help of many existing scientific artifacts, including PyTorch Paszke et al. (2019), Numpy Harris et al. (2020), huggingface, and transformers Wolf et al. (2020). We will make the Per-Pcs implementation publicly available to facilitate further research.
Appendix F Task Details
@@
Appendix G Baseline Details
@@
Appendix H Dataset Statistics
@@
Appendix I Prompt Details

### v2

@@ Abstract
Personalized large language models (LLMs) aim to tailor interactions, content, and recommendations to individual user preferences. While parameter-efficient fine-tuning (PEFT) methods excel in performance and generalization, they are costly and limit communal benefits when used individually. To this end, we introduce Personalized Pieces (Per-Pcs)111Code is available at https://github.com/TamSiuhin/Per-Pcs, a framework that allows users to safely share and assemble personalized PEFT efficiently with collaborative efforts. Per-Pcs involves selecting sharers, breaking their PEFT into pieces, and training gates for each piece. These pieces are added to a pool, from which target users can select and assemble personalized PEFT using their history data. This approach preserves privacy and enables fine-grained user modeling without excessive storage and computation demands. Experimental results show Per-Pcs outperforms non-personalized and PEFT retrieval baselines, offering performance comparable to OPPU with significantly lower resource use across six tasks. Further analysis highlights Per-Pcs’s robustness concerning sharer count and selection strategy, pieces sharing ratio, and scalability in computation time and storage space. Per-Pcs’s modularity promotes safe sharing, making LLM personalization more efficient, effective, and widely accessible through collaborative efforts.
@@
Zhaoxuan Tan, Zheyuan Liu, Meng Jiang University of Notre Dame {ztan3, mjiang2}@nd.edu
@@ 3 Experiment Settings
We adopt the Large Language Model Personalization (LaMP) benchmark Salemi et al. (2023) for our experiments, which consists of six public language model personalization tasks, including three text classification tasks (personalized citation identification, movie tagging, and producing rating) and three text generation tasks (personalized news headline generation, scholarly title generation, and tweet paraphrasing).333Task details can be found in Appendix D. We exclude the LaMP-6: Email subject generation task since it involves private data that we cannot access. We randomly select 25% of users to train the base model for task adaptation. From the remaining users, we randomly sample 100 to serve as test users for efficient and fair comparison with OPPU Tan et al. (2024). The rest of the users are used as sharer candidates who consent to share their PEFT parameters.444Statistics are presented in Table 4.
@@
We compare our proposed Per-Pcs with the non-personalized baseline, prompt-based methods (retrieval-augmented Salemi et al. (2023) and profile-augmented personalization Richardson et al. (2023)), and PEFT-based personalization methods (PEFT retrieval Zhao et al. (2024) and OPPU Tan et al. (2024)). Although PEFT retrieval has not been applied to personalization before, we employ it as a PEFT-level composition baseline. compared with Per-Pcs, OPPU requires significantly more resources, which can be seen as the upper bound for sharer personal PEFT composition. We provide more baseline details in Appendix E. For all baselines and Per-Pcs, we use Llama-2-7B Touvron et al. (2023) as the base LLM and BM25 Trotman et al. (2014) for retrieval operations to ensure efficient and fair comparisons.
@@ 5 Analysis
6 Modeling Users with Different Active Levels
7 Ablation Study
| Ablation Settings | LaMP-2 | LaMP-4 | ||
| Acc | F1 | R-1 | R-L | |
| full model | 0.410 | 0.301 | 0.193 | 0.174 |
| w/o attention | 0.340 | 0.266 | 0.176 | 0.158 |
| replace Topk-Agg. w/ Topp-Agg. | 0.390 | 0.288 | 0.169 | 0.153 |
| replace Topk-Agg. w/ Topk-Sampling | 0.383 | 0.297 | 0.172 | 0.155 |
@@ 8.2 Model Parameter Composition
Existing work has shown that performing weighted linear interpolation of model parameters leads to the composition of each model ability Li et al. (2022); Tam et al. (2023); Dou et al. (2024). This approach recycles efforts and computational resources used to create specialized models. These methods can be divided into model-, PEFT-, and piece-level compositions.
Model-level composition methods treat the entire model parameter as the minimum composition unit Wortsman et al. (2022); Zhang et al. (2024); Choshen et al. (2022); Ramé et al. (2023); Jin et al. (2022); Liu et al. (2024); Jian et al. (2024). Ilharco et al. (2022) propose the task vector, which subtracts the weights of a fine-tuned model from the pre-trained weights and conducts task vector arithmetic to enable generalization across tasks and domains. PEFT offers lightweight alternatives for fine-tuning LLMs by updating small, plug-in parameters while keeping the pre-trained weights frozen to save computational resources He et al. (2021). In PEFT-level composition, the entire PEFT module is treated as the minimum unit. By composing PEFT parameters, models can achieve task and domain generalization Shah et al. (2023); Gou et al. (2023); Zhang et al. (2023). LoRAHub Huang et al. (2023) uses a black-box optimizer to integrate specialized LoRAs, facilitating generalization to unseen tasks. Another line of work focuses on retrieving PEFT Jang et al. (2023b). LoRARetriever Zhao et al. (2024) retrieves and composes multiple LoRAs based on the given input. In piece-level composition, the minimum composition unit is a plug-in sub-component of PEFT within a specific layer. For instance, in LoRA, each low-rank update at a linear layer constitutes a "piece." Muqeeth et al. (2024) focuses on task generalization and proposes recycling PEFT pieces by employing per-token and per-piece composition under zero-shot settings.
@@ 11 Ethical Considerations
Privacy
The personalization of LLMs necessitates tailoring responses based on user-specific data, which may include sensitive or private information. The ability of an LLM to adapt its outputs to individual users raises privacy concerns, as it might inadvertently reveal personal details. This highlights the importance of implementing robust privacy safeguards in LLM personalization to ensure that personal data is handled respectfully and securely, preventing any unintended disclosures.
12 Acknowledgements
This work was supported by NSF IIS-2119531, IIS-2137396, IIS-2142827, IIS-2234058, CCF-1901059, and ONR N00014-22-1-2507.
Appendix A Computation Resources Details
@@ Appendix B Hyperparameters
Appendix C Scientific Artifacts
Appendix D Task Details
@@
Appendix E Baseline Details
@@
Appendix F Dataset Statistics
@@
Appendix G Prompt Details